Which of the following is wrong?

1.AE = C + I + G + X - M.

2.Net factor income to abroad = Factor income to abroad - Factor income from abroad.

3.GNP = NNP + depreciation

4.Net export = M - X


1
Expert's answer
2021-05-11T09:04:48-0400

Net export=M-X is wrong. Net export is the difference between the countryโ€™s value of imports and its value of exports and can be written as follows:


"Net\ export=X-M,"

here, X is exports, M is imports.

Therefore, the answer is 4. Net export = M - X.